# Interstellar medium gas heating by primordial black holes and dark matter particles

### v1

@@ III.2 PBHs and millicharged particles
The top panel of FIG. 6 shows the constraints on the PBH fraction $f_{\rm PBH}$ from interstellar medium gas heating by PBHs and millicharged particles in the Leo T dwarf galaxy. The black curve represents the upper limit of the PBH fraction $f_{\rm PBH}$ when ignoring the gas heating effects of millicharged particles, as previously reported in Ref.Takhistov2022JCAP . The remaining three curves illustrate the upper limit of $f_{\rm PBH}$ accounting for gas heating by both PBHs and millicharged particles. The red, blue, and green curves correspond to millicharged particle masses $m_{\chi}$ of $0.082$ GeV, $0.081$ GeV, and $0.80$ GeV, respectively. The excluded regions for $f_{\rm PBH}$ are indicated in the top panel of FIG. 6 as same as in FIG. 5.
@@
To clarify this, we examine the correlation between the upper bound on the charge parameter $\epsilon$ and the mass of the millicharged particle $m_{\chi}$. Since we are interested in cases where both PBHs and millicharged particles effectively contribute to the gas heating in Leo T, we consider $1\leq M_{\rm PBH}/M_{\odot}\leq 30$ and $f_{\rm PBH}=0.08$ as our benchmark case Carr2021RPP ; Takhistov2022JCAP .
The bottom panel of FIG. 6 displays the constraints on the charge parameter $\epsilon$ from interstellar medium gas heating by PBHs and millicharged particles in the Leo T dwarf galaxy. The black curve represents the upper limit of the charge parameter $\epsilon$ when ignoring the gas heating effect of PBHs, as previously reported in Ref. Wadekar2021PRD . The remaining three curves show the upper limits of $\epsilon$ when considering the gas heating contributions from both PBHs and millicharged particles. The red, blue, and green curves correspond to $M_{\rm PBH}$ = $10M_{\odot}$, $20M_{\odot}$, and $30M_{\odot}$, respectively. Excluded regions for $\epsilon$ are also indicated: SENSEI SENSEI2018PRL (SENSEI : orange), XENON Essig2012PRL ; Essig2017PRD (XENON : yellow), CSR Mahdawi2018JCAP (CSR : red), and CMB Wadekar2021PRD (CMB : gray).

### v2

@@ 3.2 PBHs and millicharged particles
The top panel of FIG. 6 shows the constraints on the PBH fraction $f_{\rm PBH}$ from interstellar medium gas heating by PBHs and millicharged particles in the Leo T dwarf galaxy. The black curve represents the upper limit of the PBH fraction $f_{\rm PBH}$ when ignoring the gas heating effects of millicharged particles, as previously reported in Ref.[26]. The remaining three curves illustrate the upper limit of $f_{\rm PBH}$ accounting for gas heating by both PBHs and millicharged particles. The red, blue, and green curves correspond to millicharged particle masses $m_{\chi}$ of $0.082$ GeV, $0.081$ GeV, and $0.80$ GeV, respectively. The excluded regions for $f_{\rm PBH}$ are indicated in the top panel of FIG. 6 as same as in FIG. 5.
@@
To clarify this, we examine the correlation between the upper bound on the charge parameter $\epsilon$ and the mass of the millicharged particle $m_{\chi}$. Since we are interested in cases where both PBHs and millicharged particles effectively contribute to the gas heating in Leo T, we consider $1\leq M_{\rm PBH}/M_{\odot}\leq 30$ and $f_{\rm PBH}=0.08$ as our benchmark case [4, 26].
The bottom panel of FIG. 6 displays the constraints on the charge parameter $\epsilon$ from interstellar medium gas heating by PBHs and millicharged particles in the Leo T dwarf galaxy. The black curve represents the upper limit of the charge parameter $\epsilon$ when ignoring the gas heating effect of PBHs, as previously reported in Ref. [28]. The remaining three curves show the upper limits of $\epsilon$ when considering the gas heating contributions from both PBHs and millicharged particles. The red, blue, and green curves correspond to $M_{\rm PBH}$ = $10M_{\odot}$, $20M_{\odot}$, and $30M_{\odot}$, respectively. Excluded regions for $\epsilon$ are also indicated: SENSEI [74] (SENSEI : orange), XENON [75, 76] (XENON : yellow), CSR [77] (CSR : red), and CMB [28] (CMB : gray).
@@ 4 Summary
Finally, we would like to comment on the heating of the gas by free-free absorption [78]. In a model that can produce a large number of low-frequency photons, these photons will be rapidly absorbed by the surrounding gas, even if the medium is only very slightly ionized. Hawking radiation from PBHs may produce many soft photons in some limits. In addition, conversions of dark photons into CMB photons (or vice versa) are more active at low frequencies, and can also produce many soft photons. These effects may give rise to the additional heating contribution, which may be important to consider. In this study, we neglect any soft photon heating. A detailed analysis of this issue is nedded in the future.